OpenStudy (dls):

Derive an expression for the G.field due to a uniform rod of Length L and mass M at a point on its perpendicular bisector at a distance d from the centre

OpenStudy (dls):



OpenStudy (dls):

\[\LARGE \frac{2GM}{\sqrt{4L^2+d^2}}\] is what I got one term seems missing

OpenStudy (yrelhan4):

again, seems like you didnt take the components right. E will be given by dE= 2dE1 sintheta where dE1=G(dm)/(d^2 + x^2) where dm=m/l dx sintheta=d/(d^2 + x^2) integrate with respect to x from 0 to l/2